OpenStudy (anonymous):

Alice purchased a house using a fixed rate mortgage. The annual interest rate is 5.6% compounded monthly for 30 years. The mortgage amount is $205,000. What is the monthly payment on the mortgage? $601.33 $1,176.86 $1,189.00 $1,271.15

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@TuringTest

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@golfergirl93

OpenStudy (mertsj):

\[M=P\frac{i(1+i)^n}{(1+i)^n-1}\]

OpenStudy (mertsj):

P= principal i = interest rate n = number of payments

OpenStudy (anonymous):

and what would that give you cause m not getting the correct answer

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

@zimmah

OpenStudy (anonymous):

@Zale101

OpenStudy (anonymous):

@charlotte123

OpenStudy (anonymous):

@wolfe8

OpenStudy (anonymous):

I know I am a tad late. But for anyone else who may need it \[M = P [ i(1 + i)^n ] / [ (1 + i)^n - 1]\] Would be the formula like @mertsj said. I believe when you said you got the wrong answer, you did not divide the interest rate by 12 since you were given the annual interest. Here is my work: P = 205,000 i = .056/12 = 0.00466666666 (since you were given the annual interest and we need the monthly payment) n = 30*12 = 360 (since you were given the years and you need the months) M = 205,000 [ 0.00466666666(1 + 0.00466666666)^360 ] / [ (1 + 0.00466666666)^360 - 1] M = 205,000 [ 0.00466666666(1.00466666666)^360 ] / [ (1.00466666666)^360 - 1] M = 205,000 [ 0.00466666666(5.34462910282) ] / [ (5.34462910282) - 1] M = 205,000 [.02494160244] / [4.34462910282] M = 205,000 [.0057407898] M = $1,176.86